What is the exact value of cos210degrees sin^2 60degrees/cos(-135degrees)
Answer by stanbon(75887)

What is the exact value of cos210degrees sin^2 60degrees/cos(-135degrees)
------------------
cos(210) = (-sqrt3)/2
---------------
sin60 = (sqrt3)/2 ; so sin^2(60) = 3/4
------------------
cos(-135) = (-sqrt2)/2
==========================
So cos210degrees sin^2 60degrees/cos(-135degrees)
= [(-sqrt3)/2][3/4]/[(-sqrt2)/2
=(3/4)[sqrt(3)/2]
= [3sqrt(3)]/[4sqrt2]
= [3sqrt(6)]/8
= (3/8)sqrt(6)
===================
Cheers,
Stan H.